OpenStudy (anonymous):

-2x^2y, if x = -3 and y = -1 -12 -18 18 12

OpenStudy (anonymous):

so to solve this you would put the numbers in place of Y and X \[-2\times-3^2\times-1=\]

OpenStudy (anonymous):

the equation messed up D:

OpenStudy (anonymous):

Evaluate -2x^2y, if x = -3 and y = -1 is what it says

OpenStudy (anonymous):

but for x instead of x plug in -3, and instead of y, plug in -1 :P

OpenStudy (anonymous):

that sounded like a foreign language to me haha

OpenStudy (anonymous):

idk what the ^ thingie is so you're on your own there but just solve the equation and you should get it

OpenStudy (anonymous):

to the second power haha

OpenStudy (anonymous):

ah well it's really wuite easy to solve, just plug in the numbers :P

OpenStudy (anonymous):

is it 12?